OpenStudy (carolina4567):

Which of the following equations have the same solution? Solve all three equations separately, and explain each step. I. 3x + 2 = 5x – 6 II. 2x – 2 = 8 III. the quantity of x plus 1, divided by 5 equals the quantity of 2x minus 2, divided by 6

OpenStudy (carolina4567):

I need a little help :)

OpenStudy (cwrw238):

rearrange the first equation to the same form as the second and see what you get

OpenStudy (carolina4567):

so would it be 3x - 5x = 2 + 6

OpenStudy (invinceble):

have no clue

OpenStudy (anonymous):

@invinceble if you have no clue it is best not to say it

OpenStudy (carolina4567):

is that correct @cwrw238

OpenStudy (cwrw238):

no

OpenStudy (cwrw238):

subtract 3x from both sides then add 6 to both sides

OpenStudy (anonymous):

@cwrw238 is right. You need to subract from both sides

OpenStudy (anonymous):

subtract*

OpenStudy (carolina4567):

Hmm so that would be like 3x + 2 = 5x – 6 -3 -3 Is that ok ?

OpenStudy (cwrw238):

no subtract 3x not 3

OpenStudy (anonymous):

Math can be so confusing sometimes XD

OpenStudy (carolina4567):

Oh 3x + 2 = 5x – 6 -3x -3x

OpenStudy (carolina4567):

yes it can @PaulaLovesSchool13

OpenStudy (cwrw238):

3x + 2 = 5x - 6 - 3x -3x 2 = 2x - 6 +6 + 6 8 = 2x what what is x?

OpenStudy (carolina4567):

x = 4

OpenStudy (anonymous):

Good!

OpenStudy (cwrw238):

ok now solve the second equation

OpenStudy (carolina4567):

2x – 2 = 8 + 2 +2 2x = 10 x = 5 Correct ?

OpenStudy (cwrw238):

right

OpenStudy (anonymous):

Yes, that should be right. Good job Carolina :)

OpenStudy (carolina4567):

Yay

OpenStudy (anonymous):

XD

OpenStudy (carolina4567):

Ok last one is rlly hard

OpenStudy (anonymous):

Uh oh...

OpenStudy (anonymous):

That is hard...ugh. Hold one sec

OpenStudy (cwrw238):

with this one you get rid of the fractions first - one way is to cross multiply

OpenStudy (anonymous):

The mathlete has some nice tips. Medal for you!

OpenStudy (cwrw238):